Southwest Sands currently has 18,500 shares of stock outstanding. It is considering issuing $107,000 of debt at an interest rate of 6.8 percent. The break-even level of EBIT between these two capital structure options is $67,000. How many shares of stock will be repurchased if the company undergoes the recapitalization? Ignore taxes. Multiple Choice 1,578.54 shares 2,176.47 shares 2,009.04 shares 1,908.59 shares 1,722.04 shares

PLEASE SHOW WORK
Answer #1

Debt issued for $107,000

Interest rate = 6.80%

Interest expenses = 107000*6.80%= $7,276

Shares before repurchase= 18500

Breakeven level of EBIT = 67000

Breakeven EBIT is that point where EPS under two capital structure is same. Equation is below:

EBIT/No. of shares before repurchase =( EBIT - Interest)/No. of shares after repurchase

67000/18500 = (67000-7276)/No. of shares after repurchase

3.621621622 (67000-7276)/No. of shares after repurchase

No. of shares after repurchase= (67000-7276)/ 3.621621622 =16,490.96

No. of shares repurchased =No of shares before repurchases - No. of shares after repurchases

18500- 16,490.96

=2,009.04

So, no. of shares repurchased are 2009.04

Please thumbs up.